NEW LABORATORY MEASUREMENTS OF RHOMBOIDAL SiC_{3}

\underline{CARL A. GOTTLIEB} and PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138.

Rhomboidal SiC_{3}, the highly polar planar ring with C_{2v} symmetry and a transannular C—C bond, was detected in our laboratory about 10 years ago,^{a} and soon afterwards was identified with a radio telescope in the expanding envelope of IRC+10216.^{b} Recently a sensitive spectral line survey of IRC+10216 was made with the Submillimeter Array (SMA) in the 300 − 355 GHz range with a 3′′ × 2′′ synthesized beam. Many new lines were detected in this survey. Most are from high rotational transitions of molecules that are known in IRC+10216, but some of the lines are quite narrow and more than 10 of these are unassigned.^{c} In support of the SMA observations we have extended the earlier laboratory measurements by Apponi et al. from 286 GHz and K_{a} ≤ 6, to 450 GHz and K_{a} ≤ 20 from rotational levels as high as 825 K above ground. As a result uncertainties in the predicted spectrum for lines with high K_{a} have been reduced by as much as two orders of magnitude, which should aid the assignment of SiC_{3} in the SMA survey and in future observations with ALMA.

aA. J. Apponi, M. C. McCarthy, C. A. Gottlieb, and P. Thaddeus, Journ. Chem. Phys. 111, 3911 (1999).

bA. J. Apponi, M. C. McCarthy, C. A. Gottlieb, and P. Thaddeus, Astrophys. Journ. Lett. 516, L103 (1999).

cN. A. Patel, K. H. Young, S. Brünken, R. W. Wilson, P. Thaddeus, K. M. Menten, M. Reid, M. C. McCarthy, Dinh-V Trung, C. A. Gottlieb, and A. Hedden, Astrophys. Journ., in press (2009).